OpenStudy (anonymous):

can you help me solve each equation system by the addition method 4x+2y=6 4x-5y=-1.

OpenStudy (anonymous):

anythingwill help lol

OpenStudy (anonymous):

multiply 2nd equation by -1

OpenStudy (anonymous):

substitute y=1 into the equation4x+2y=6 4x+2(1)=6 4x=6−2 4x=4 x=?

OpenStudy (anonymous):

so x=1?

OpenStudy (anonymous):

:) mhmm

OpenStudy (anonymous):

would I do the same thing for y?

OpenStudy (anonymous):

•/| 4*x+2*y-6 = 0| 4*x-(5*y)+1 = 0 • •We try to solve the equation: 4*x+2*y-6 = 0 •4*x+2*y-6 = 0 // - 4*x-6 •2*y = -(4*x-6) // : 2 •y = (-(4*x-6))/2 •y = (6-(4*x))/2 • •We insert the solution into one of the initial equations of our system of equations •We get a system of equations: •/| 4*x-(5*((6-(4*x))/2))+1 = 0| y = (6-(4*x))/2 •(-1*5*(6-4*x))/2+4*x+1 = 0 •14*x-15+1 = 0 •14*x-14 = 0 • •14*x-14 = 0 // + 14 •14*x = 14 // : 14 •x = 14/14 •x = 1 • •We insert the solution into one of the initial equations of our system of equations •For y = (6-(4*x))/2: • •y = (6-(1*4))/2 •y = (6-4)/2 •y = 2/2 •y = 1 • •We get a system of equations: •/| y = 1| x = 1

OpenStudy (anonymous):

so ur y and x are both equal to 1 :)

OpenStudy (anonymous):

I get it now!! Thank you :)

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (anonymous):

If Your Need Anything ... everyone is here to help noah don't be shy :)